OpenStudy (anonymous):

Solve the system by using the substitution method 2x-3y=3 -2x+y=-4

OpenStudy (mertsj):

It would be easiest to just add the two equations which will get rid of the x. Then you can solve for y

OpenStudy (across):

Inopeki should help you with this one.

OpenStudy (mertsj):

If you must use substitution, it would be easiest to solve the second equation for y

OpenStudy (anonymous):

what would y be?

OpenStudy (mertsj):

What did you get when you added 2x to both sides?

OpenStudy (anonymous):

get one variable by itself in one of the equations, then sub into the other equation (1) 2x-3y=3 (2) -2x+y=-4 from (2) y=-4+2x sub in (1) 2x-3(-4+2x)=3 solve for x, then plug in what you got for x into one of the original equations and solve for y does that make sense?

OpenStudy (anonymous):

yes! thank you!

OpenStudy (anonymous):

your welcome!

OpenStudy (anonymous):

do you know how to solve for x?

OpenStudy (anonymous):

i didnt get the answer..i got x=-9/-4

OpenStudy (anonymous):

did the question ask for what x is or y is?

OpenStudy (anonymous):

both

OpenStudy (anonymous):

idk how to get em

OpenStudy (anonymous):

oh okay, does your question sheet have the answers on it?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

okay so you have x=-9/-4, the negatives cancel out so you have x=9/4 now sub 9/4 into one of the original equations for example: use equation 2 -2x+y=-4 -2(9/4)+y=-4 -9/2+y=-4 y=-4+9/2 y=17/2

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

does that make sense? do you understand how i got that answer?